# Python-Based Quantum Chemistry Calculations with GPU Acceleration

### v1

@@ Abstract
To meet the increasing demand of quantum chemistry calculations in data-driven chemical research, the collaboration between industrial stakeholders and the quantum chemistry community has led to the development of GPU4PySCF, a GPU-accelerated Python package. This open-source project is accessible via its public GitHub repository at https://github.com/pyscf/gpu4pyscf. This paper outlines the primary features, innovations, and advantages of this package. When performing Density Functional Theory (DFT) calculations on modern GPU platforms, GPU4PySCF delivers 30 times speedup over a 32-core CPU node, resulting in approximately 90% cost savings for most DFT tasks. The performance advantages and productivity improvements have been found in multiple industrial applications, such as generating potential energy surfaces, analyzing molecular properties, calculating solvation free energy, identifying chemical reactions in lithium-ion batteries, and accelerating neural-network methods. To make the package easy to extend and integrate with other Python packages, it is designed with PySCF-compatible interfaces and Pythonic implementations. This design choice enhances its coordination with the Python ecosystem.
@@ 1 Introduction
In this work, we have developed a Python-based and GPU-optimized framework, building upon the foundations of PySCF. This initiative aims to advance quantum chemistry calculations on GPUs by focusing on several key areas: 1) Prioritizing density fitting algorithms, which are inherently more GPU-friendly, while offering limited support for direct Self-Consistent Field (SCF) methods. 2) Adapting our codebase to leverage modern GPU architectures, such as tensor cores, to achieve up to twice the speed in tensor contractions. 3) Integrating a wide array of open-source packages from the quantum chemistry community within our GPU4PySCF package, thus supporting sophisticated DFT methods, popular basis sets, solvent models, chemical property calculations, geometry optimization, and transition state search methodologies. This comprehensive approach signifies a notable advancement in the field of quantum chemistry, bridging the gap between traditional computational methods and the cutting-edge capabilities offered by modern GPU technology.
We underscore the notable advantages of GPU4PySCF as follows:
  • Speed: GPU4PySCF provides performance that is equivalent to that of 600-1000 CPU cores for running conventional CPU-based quantum chemistry software. As a result, it offers substantial cost savings, for instance, up to 90% when utilizing the NVIDIA A100-80G cards.
@@
  • Open Source: GPU4PySCF benefits from the support of the PySCF community, the largest open-source quantum chemistry community, fostering collaborative development and innovation.
  • Application Driven: Specifically optimized and tailored for industrial applications, GPU4PySCF excels in tasks such as DFT calculations, offering practical advantages for commercial research and development.
Detailed performance metrics and benchmarks are provided in Section 2. These calculations have been rigorously cross-validated with Q-Chem 6.1, ensuring reliability and accuracy. Furthermore, Section 3 briefly illustrates how GPU4PySCF can be utilized to navigate the potential energy landscape, analyze quantum features, compute thermodynamic properties, estimate solvation free energies, quantify chemical reactions, and integrate with neural network models. This comprehensive approach showcases the framework’s versatility and potential to revolutionize quantum chemistry computations through GPU acceleration.
@@ 2.1 Dataset and Performance
We focus on the benchmarks with density fitting scheme, which is efficient for small molecules with less than 200 atoms. A small and diverse dataset is constructed for the benchmarking the fundamental modules in GPU4PySCF. We select small molecules ($<$ 100 atoms) from the supplement information in [20] and two medium-sized molecules from the supplement information in [8]. The newly constructed dataset includes 8 elements (H, C, O, N, Mg, S, Cl, P), and 13 small molecules with 20-168 atoms. The details of the dataset are provided in Appendix A. The xyz files and benchmarking details can be found on GitHub. Other transition metals, such as Nb, Ti, Ru and so on, are also supported in GPU4PySCF, although they are not included in this dataset. This capability will be shown in Sec. 3.4.1. Larger molecules with more than 168 atoms still can be calculated with the direct SCF scheme in GPU4PySCF. The detailed performance of direct SCF scheme will be disclosed in the future work. The availability of direct SCF scheme can be found in Sec. 4.2.
For a typical DFT protocol with def2-TZVPP, B3LYP, and (99,590) grids, we time SCF, gradient, and Hessian calculations on NVIDIA A100-80G. For the smallest molecule in the dataset, SCF, gradient, and Hessian take 1.7, 0.68, and 43 seconds respectively. For the largest molecule in the dataset, SCF, gradient, and Hessian take 309 seconds, 58 seconds, and 21 hours respectively. Hessian calculation is the most expensive module of the vibrational analysis. With GPU4PySCF, one can perform the vibrational analysis for large molecules with 168 atoms in one day. The vibrational analysis of moderately sized (100-200 atoms) molecules is prohibitively expensive with traditional quantum chemistry softwares [21]. The time can be further reduced with the recent GPUs such as H100. The time scaling curve of SCF, gradient, and Hessian calculations is shown in Figure 1. Theoretically, SCF and gradient calculations are quartic scaling with respect to the system size, while Hessian calculation is $O(N^{5})$. The slope of the scaling curve decreases as the system size gets larger, due to sparsity. The bottleneck of calculating a large molecule is the CPU memory storage. A CPU-GPU hybrid strategy is employed to store GTO integrals in this work. For small molecules, all the intermediate variables are stored in GPU memory. For the large system, the variables are stored in CPU memory. The data in CPU memory is asynchronously transferred to GPU memory for consumption. The performance does not drop significantly in the scaling curve.
@@ 2.3 Speedup of DFT with Implicit Solvent Models
Vitamin C 41.1 28.4 106.4 41.6 27.3 104.9
Inosine 65.5 38.9 147.4 65.6 40.2 146.2
Bisphenol A 76.2 43.3 162.7 76.7 45.4 161.3
Mg Porphin 84.0 49.0 162.2 87.3 49.8 160.1
Penicillin V 83.2 44.1 170.5 84.0 43.4 169.4
@@ 3 Applications
In this section, we showcase the capabilities of GPU4PySCF with several industrial applications. A complete list of capabilities will be shown in Section 4.1.
@@ 3.1.3 Lithium Ion Solvation Structure
Most of the clusters we examined have a coordination number of 4, consisting of either solvent or anion molecules, or a combination of the two. Their binding energies range from -0.065 eV to -1.078 eV, indicating significant differences in the energy stability of the clusters. We observe that clusters composed solely of Li+ and FSI- exhibit the most positive binding energies compared to those containing solvents. This observation aligns with chemical intuition, suggesting that the salt dissolution of LiFSI in an EC-DMC polar solvent mixture is preferred, manifesting the fundamental requirement of solvent functionality in Li-battery liquid electrolyte applications. Besides, clusters with the same composition can exhibit different energies if their structures vary, as shown in Figure 6 for the $\text{Li}^{+}(\text{DMC})(\text{EC})(\text{FSI}^{-})_{2}$ cluster, suggesting adequate geometric sampling for DFT calculations is necessary for accurate representation of solvation structure stabilities, emphasizing the importance of computational efficiency provided in our code.
@@ 3.2.1 CHELPG Charge Calculation
Followed by the work of Herbert and co-workers [56], CHELPG is implemented in our package, and the computational efficiency has been enhanced with the aid of GPUs. Our implementation uses the Cartesian grid with the default spacing 0.3 Å and the radial extent of the CHELPG grid 2.8 Å as recommended in Ref. [54]. Besides, the smoothing function used in Ref. [56] is also implemented, where the Bondi radius is used [57] (the radius of hydrogen is changed to 1.1 Å). Consistent results can be obtained when employing identical settings with Q-Chem. As shown in Figure 7, in GPU4PySCF, the calculation speed of CHELPG has been greatly improved approximately 10 times versus CPU-based quantum chemistry software [5]. This improvement in computational efficiency is due to the efficient implementation of molecular integrations (calculating the electrostatic potential at grids) on GPU. All the calculations in this subsection are performed using B3LYP functional and def2-SVPD basis with density fitting used.
@@ 3.3 Solvation Free Energy
Implicit solvent models are computational tools designed to efficiently simulate the effect of a solvent on molecular systems, emphasizing the solvent’s overall impact rather than detailing every interaction between solvent and solute molecules. In the realm of quantum chemistry, electrostatic interactions between solvent and solute molecules are captured using the Polarizable Continuum Model (PCM), which offers a nuanced understanding of molecular properties in solution. In GPU4PySCF, users can access four variants of PCM models: C-PCM, COSMO, IEF-PCM, and SS(V)PE. For accurate prediction of solvation energies, accounting for non-electrostatic interactions such as cavitation, Pauli repulsion, dispersion, and hydrogen bonding is essential. These contributions are encompassed within more advanced models like SMD [26], CMIRS [65], and COSMO-RS [66] models, with the SMD model being particularly prevalent in quantum chemistry calculations and available from GPU4PySCF.
The implementation of the SMD model varies across different quantum chemistry software platforms. And the algorithm has been updated with minor modifications over time. We benchmark our implementation against Q-Chem and Gaussian 03, and we note several unique aspects in our approach:
@@
  • Atomic SASA and molecular SASA are calculated numerically using Lebedev quadrature, diverging from the analytical formulation found in [68].
@@
PySCF/M06-2X/TZ 0.69 0.68 5.9 4.2 2.4 5.6
PySCF/M06-2X/6-31G* 0.71 0.65 5.9 4.6 2.1 4.8
PySCF/B3LYP/TZ 0.79 0.69 5.7 3.8 3.1 6.3
PySCF/B3LYP/6-31G* 0.89 0.70 5.8 4.3 2.5 5.4
PySCF/HF/6-31G* 0.84 0.74 6.2 5.4 3.1 3.8
@@
Table 3: Mean unsigned error (MUE) in solvation free energies (kcal/mol) with different protocols. The detailed protocols for G03 are described in the original SMD paper [26], Table 13. The data for Q-Chem is take from [71]. In all calculations, the bulk electrostatic contribution is calculated with IEF-PCM. Density fitting is applied for all PySCF calculations. 2346 solvation free energies for neutrals and 332 solvation free energies for ions are used for G03. 2369 solvation free energies and 361 solvation free energies are used for PySCF results. The uncertainty of the reference data are $\pm$ 0.2 kcal/mol for neutral solutes and $\pm$ 3 kcal/mol for ions. ‘TZ’ is short for def2-TZVPP. ‘NAQ’ is short for nonaqueous.
@@ 3.4.1 Transition State Search
We benchmark the results with $\omega$B97M-V, which is one of the best performer in MOBH35 database [76]. In the dataset, the reference energies are calculated with $\mathrm{CCSD(T)/CBS_{w1}+\Delta(T)/TZVPP}$. The barrier heights and reaction energies are presented in Fig. 9. The average absolute deviation (MAD) of the reaction energy we calculated in the subset of MOBH35 is 1.75 kcal/mol, which is similar to MAD (1.7 kcal/mol) of the entire database reported in [76]. The re-optimized transition states are almost identical to the geometries reported in [76]. The difference of the reaction energy between the original geometry and re-optimized geometries is 0.01 kcal/mol in MAD. That verifies the robustness of transition state optimization scheme.
@@ 3.4.2 Electrochemical Stability Windows of Electrolytes
In our study, we calculated the redox potentials of some molecules reported in Ref. [80], and compared our results with those previously reported in the same reference. It should be noted that our calculations were performed using a thermodynamic cycle, with computational parameters that slightly differ from those used in Ref. [80]. As shown in Figure 10, despite differences in the computational approach and parameters from those described in the original literature, the results obtained are remarkably similar. Consequently, we posit that our package possesses the capability to calculate redox potentials, thereby enabling the prediction of electrochemical windows. Considering the computational efficiency, oxidation and reduction potential calculations involves the optimization of geometric structures in vacuum and solvent, as well as the calculation of Hessian in vacuum. The computational acceleration using GPU4PySCF with respect to Q-Chem for these processes has been discussed in previous sections. Consequently, the acceleration ratio for the calculation of oxidation-reduction potential is determined by the lowest acceleration ratio among these individual parts.
@@ 3.5 Incorporating with Neural Networks
4 Community and Ecosystem
PySCF has been a constantly growing open-source community. It has been integrated into various open-source packages as the interface to quantum chemistry. For example,
  • Quantum Computing. For instance, Google’s quantumlib, IBM’s Qiskit, NVIDIA’s CUDA Quantum, Tencent’s TenCirChem.
  • Deep Learning. For instance, DeepMind’s FermiNet [92] and DM21, DeepModeling’s DeePKS-kit [97], ByteDance’s DeepSolid and LapNet.
  • Quantum Chemistry Workflows. For instance, Microsoft’s AiiDA.
Without changing too many code, those works can be accelerated with GPU4PySCF.
4.1 Compatibility with PySCF
language = Python, keywordstyle = [2], otherkeywords = to_gpu, morekeywords = [2]to_gpu, GPU4PySCF objects are designed to be compatible with the corresponding PySCF objects. A PySCF object can be converted into a GPU4PySCF object with calling to_gpu function when the functionality is implemented on GPU. The kernel execution will be accelerated with GPU. On the other hand, a GPU4PySCF object can be converted into a PySCF object with to_gpu. Then PySCF features if not accelerated still can be applied. Here is an example

### v2

# Enhancing GPU-acceleration in the Python-based Simulations of Chemistry Framework

@@ Abstract
We describe our contribution as industrial stakeholders to the existing open-source GPU4PySCF project (https://github.com/pyscf/gpu4pyscf), a GPU-accelerated Python quantum chemistry package. We have integrated GPU acceleration into other PySCF functionality including Density Functional Theory (DFT), geometry optimization, frequency analysis, solvent models, and density fitting technique. Through these contributions, GPU4PySCF v1.0 can now be regarded as a fully functional and industrially relevant platform which we demonstrate in this work through a range of tests. When performing DFT calculations on modern GPU platforms, GPU4PySCF delivers 30 times speedup over a 32-core CPU node, resulting in approximately 90% cost savings for most DFT tasks. The performance advantages and productivity improvements have been found in multiple industrial applications, such as generating potential energy surfaces, analyzing molecular properties, calculating solvation free energy, identifying chemical reactions in lithium-ion batteries, and accelerating neural-network methods. With the improved design that makes it easy to integrate with the Python and PySCF ecosystem, GPU4PySCF is natural choice that we can now recommend for many industrial quantum chemistry applications.
@@ 1 Introduction
In this work, we will present our contributions to the GPU4PySCF package and related modules to enhance its functionality. GPU4PySCF is an open-source project that was initiated in the Chan group [20] to add GPU acceleration to key parts of the PySCF computational workflow. At the time that we started our work, GPU4PySCF contained CUDA kernels for 4-center two-electron integrals and GPU accelerated direct SCF (Hartree-Fock) functionality [20]. Our contributions enhance the capabilities of this module in several key areas: 1) through density fitting algorithms. These adapt the Rys quadrature CUDA kernels in GPU4PySCF to the 3-center quantities for density fitting. However, once the density fitted integrals are computed, the remaining operations are tensor operations, which are inherently more GPU-friendly. The density fitting implementation is optimized for small to medium sized molecules, and thus complements the existing functionality for direct Self-Consistent Field (SCF) methods that can be extended to larger molecules. 2) Optimizing our codebase to leverage modern GPU architectures, such as tensor cores, to achieve up to twice the speed in tensor contractions. 3) Integrating a wide array of open-source packages from the quantum chemistry community into the PySCF package, thus supporting sophisticated DFT methods, popular basis sets, solvent models, chemical property calculations, geometry optimization, and transition state search methodologies. Through these comprehensive contributions, we can now consider the GPU4PySCF package to be feature-rich and an industrially relevant quantum chemistry tool.
With our contributions, we underscore the notable advantages of GPU4PySCF as follows:
  • Speed: GPU4PySCF now provides performance that is equivalent to that of 600-1000 CPU cores for running conventional CPU-based quantum chemistry software. As a result, it offers substantial cost savings, for instance, up to 90% when utilizing the NVIDIA A100-80G cards.
@@
  • Open Source: Our contributions to GPU4PySCF benefits from the support of the PySCF community, the largest open-source quantum chemistry community, fostering collaborative development and innovation.
  • Application Driven: With our contributions that are specifically optimized and tailored for industrial applications, GPU4PySCF excels in tasks such as DFT calculations, offering practical advantages for commercial research and development.
Detailed performance metrics and benchmarks are provided in Section 2. These calculations have been rigorously cross-validated with Q-Chem 6.1, ensuring reliability and accuracy. Furthermore, in Section 3 briefly illustrates how GPU4PySCF v1.0 can be utilized to navigate the potential energy landscape, analyze quantum features, compute thermodynamic properties, estimate solvation free energies, quantify chemical reactions, and integrate with neural network models. This comprehensive approach showcases the framework’s versatility and potential to revolutionize quantum chemistry computations through GPU acceleration.
@@ 2.1 Dataset and Performance
We focus on benchmarks with the density fitting scheme, using the optimized implementation that we have contributed. This is efficient for small molecules with less than 200 atoms. A small and diverse dataset is constructed for the benchmarking the fundamental modules in GPU4PySCF. We select small molecules ($<$ 100 atoms) from the supplement information in [21] and two medium-sized molecules from the supplement information in [8]. The newly constructed dataset includes 8 elements (H, C, O, N, Mg, S, Cl, P), and 13 small molecules with 20-168 atoms. The details of the dataset are provided in Appendix A. The xyz files and benchmarking details can be found on GitHub. Other transition metals, such as Nb, Ti, Ru and so on, are also supported in GPU4PySCF, although they are not included in this dataset. This capability will be shown in Sec. 3.4.1. Larger molecules with more than 168 atoms can be calculated within the existing direct SCF scheme in GPU4PySCF [20], which we do not focus on here.
For a typical DFT protocol with def2-TZVPP, B3LYP, and (99,590) grids, we record the wall clock time for SCF, gradient, and Hessian calculations on NVIDIA A100-80G. For the smallest molecule (Vitamin C) in the dataset, SCF, gradient, and Hessian take 1.7, 0.68, and 43 seconds respectively. For the largest molecule (Valinomycin) in the dataset, SCF and gradient calculations take 309 seconds and 58 seconds respectively. The Hessian calculation of Sphingomyelin (84 atoms) takes 30 minutes. Hessian calculation is the most expensive module of the vibrational analysis. The vibrational analysis of moderately sized (100-200 atoms) molecules is prohibitively expensive with traditional quantum chemistry softwares [22]. Unfortunately, coupled-perturbed SCF iterations for (Azadirachtin, Taxol, Valinomycin) does not converge. The time scaling curve of SCF, gradient, and Hessian calculations is shown in Figure 1. Theoretically, SCF and gradient calculations are quartic scaling with respect to the system size, while Hessian calculation is $O(N^{5})$. The slope of the scaling curve decreases as the system size gets larger, due to sparsity. The bottleneck of calculating a large molecule is the CPU memory storage. A CPU-GPU hybrid strategy is employed to store GTO integrals in this work. For small molecules, all the intermediate variables are stored in GPU memory. For the large system, the variables are stored in CPU memory. The data in CPU memory is asynchronously transferred to GPU memory for consumption.
@@ 2.3 Speedup of DFT with Implicit Solvent Models
Vitamin C 39.8 21.6 107.2 31.0 18.3 100.1
Inosine 64.3 30.4 162.3 54.3 30.6 158.2
Bisphenol A 72.8 35.9 158.7 67.1 37.6 155.2
Mg Porphin 86.8 40.9 168.5 81.1 41.4 166.4
Penicillin V 78.0 36.8 186.0 81.0 36.7 186.5
@@ 3 Applications
In this section, we take the opportunity to showcase the capabilities of of GPU4PySCF v1.0 with our new contributions in several industrial applications. A complete list of capabilities will be shown in Section 4.2.
@@ 3.1.3 Lithium Ion Solvation Structure
Most of the clusters we examined have a coordination number of 4, consisting of either solvent or anion molecules, or a combination of the two. Their binding energies range from -0.065 eV to -1.078 eV, indicating significant differences in the energy stability of the clusters. We observe that clusters composed solely of Li+ and FSI- exhibit the most positive binding energies compared to those containing solvents. This observation aligns with chemical intuition, suggesting that the salt dissolution of LiFSI in an EC-DMC polar solvent mixture is preferred, manifesting the fundamental requirement of solvent functionality in Li-battery liquid electrolyte applications. Besides, clusters with the same composition can exhibit different energies if their structures vary, as shown in Figure 6 for the $\text{Li}^{+}(\text{DMC})(\text{EC})(\text{FSI}^{-})_{2}$ cluster, suggesting adequate geometric sampling for DFT calculations is necessary for accurate representation of solvation structure stabilities, emphasizing the importance of computational efficiency provided in the code.
@@ 3.2.1 CHELPG Charge Calculation
Followed by the work of Herbert and co-workers [57], CHELPG is implemented in the package, and the computational efficiency has been enhanced with the aid of GPUs. Our implementation uses the Cartesian grid with the default spacing 0.3 Å and the radial extent of the CHELPG grid 2.8 Å as recommended in Ref. [55]. Besides, the smoothing function used in Ref. [57] is also implemented, where the Bondi radius is used [58] (the radius of hydrogen is changed to 1.1 Å). Consistent results can be obtained when employing identical settings with Q-Chem. As shown in Figure 7, in GPU4PySCF, the calculation speed of CHELPG has been greatly improved approximately 10 times versus CPU-based quantum chemistry software [5]. This improvement in computational efficiency is due to the efficient implementation of molecular integrations (calculating the electrostatic potential at grids) on GPU. All the calculations in this subsection are performed using B3LYP functional and def2-SVPD basis with density fitting used.
@@ 3.3 Solvation Free Energy
Implicit solvent models are computational tools designed to efficiently simulate the effect of a solvent on molecular systems, emphasizing the solvent’s overall impact rather than detailing every interaction between solvent and solute molecules. In the realm of quantum chemistry, electrostatic interactions between solvent and solute molecules are captured using the Polarizable Continuum Model (PCM), which offers a nuanced understanding of molecular properties in solution. In our enhancement of GPU4PySCF, users can access four variants of PCM models: C-PCM, COSMO, IEF-PCM, and SS(V)PE. For accurate prediction of solvation energies, accounting for non-electrostatic interactions such as cavitation, Pauli repulsion, dispersion, and hydrogen bonding is essential. These contributions are encompassed within more advanced models like SMD [27], CMIRS [66], and COSMO-RS [67] models, with the SMD model being particularly prevalent in quantum chemistry calculations and available from GPU4PySCF.
The implementation of the SMD model varies slightly across different quantum chemistry software platforms. And the algorithm has been updated with minor modifications over time. Compared with other implementations in Q-Chem and G03, we point out several possible differences in our implementation:
@@
  • Atomic SASA, molecular SASA, and CDS contributions are calculated using the Fortran code from NWChem [69]. Those calculations are also re-implemented in Python, where atomic SASA and molecular SASA are calculated numerically using Lebedev quadrature, diverging from the analytical formulation in [70]. The difference is usually negligible for calculating the CDS contribution.
@@
GPU4PySCF/M06-2X/TZ 0.72 0.68 6.0 4.2 2.4 5.5
GPU4PySCF/M06-2X/6-31G* 0.74 0.65 5.9 4.6 2.1 4.8
GPU4PySCF/B3LYP/TZ 0.82 0.68 5.7 3.8 3.1 6.2
GPU4PySCF/B3LYP/6-31G* 0.94 0.70 5.8 4.3 2.5 5.4
GPU4PySCF/HF/6-31G* 0.91 0.74 6.2 5.4 3.1 3.8
@@
Table 3: Mean unsigned error (MUE) in solvation free energies (kcal/mol) with different protocols. The detailed protocols for G03 are described in the original SMD paper [27], Table 13. The data for Q-Chem is take from [73]. In all calculations, the bulk electrostatic contribution is calculated with IEF-PCM. Density fitting is applied for all GPU4PySCF calculations. 2346 solvation free energies for neutrals and 332 solvation free energies for ions are used for G03 [27] 2368 solvation free energies for neutrals and 362 solvation free energies for ions are used for PySCF results. The uncertainty of the reference data are $\pm$ 0.2 kcal/mol for neutral solutes and $\pm$ 3 kcal/mol for ions. ‘TZ’ is short for def2-TZVPP. ‘NAQ’ is short for nonaqueous.
@@ 3.4.1 Transition State Search
We benchmark the results with $\omega$B97M-V, which is one of the best performer in MOBH35 database [78]. In the dataset, the reference energies are calculated with $\mathrm{CCSD(T)/CBS_{w1}}$ + $\mathrm{\Delta(T)/TZVPP}$. The barrier heights and reaction energies are presented in Fig. 9. The average absolute deviation (MAD) of the reaction energy we calculated in the subset of MOBH35 is 1.75 kcal/mol, which is similar to MAD (1.7 kcal/mol) of the entire database reported in [78]. The re-optimized transition states are almost identical to the geometries reported in [78]. The difference of the reaction energy between the original geometry and re-optimized geometries is 0.01 kcal/mol in MAD. That verifies the robustness of transition state optimization scheme.
@@ 3.4.2 Electrochemical Stability Windows of Electrolytes
In our study, we calculated the redox potentials of some molecules reported in Ref. [82], and compared our results with those previously reported in the same reference. It should be noted that our calculations were performed using a thermodynamic cycle, with computational parameters that slightly differ from those used in Ref. [82]. As shown in Figure 10, despite differences in the computational approach and parameters from those described in the original literature, the results obtained are remarkably similar. Consequently, we posit that the package possesses the capability to calculate redox potentials, thereby enabling the prediction of electrochemical windows. Considering the computational efficiency, oxidation and reduction potential calculations involves the optimization of geometric structures in vacuum and solvent, as well as the calculation of Hessian in vacuum. The computational acceleration using GPU4PySCF with respect to Q-Chem for these processes has been discussed in previous sections. Consequently, the acceleration ratio for the calculation of oxidation-reduction potential is determined by the lowest acceleration ratio among these individual parts.
@@ 3.5 Incorporating with Neural Networks
4 Compatibility and Functionalities
As an open-source Python library, PySCF has been widely adopted in quantum computing, deep learning, and various quantum chemistry workflows, most of which extensively utilize GPU acceleration. In this section, we introduce improvements to the interfaces and outline the functionalities that have been accelerated. With minimal code changes, these tasks can be sped up using GPU4PySCF v1.0.
4.1 Improved Compatibility with PySCF
language = Python, keywordstyle = [2], otherkeywords = to_gpu, morekeywords = [2]to_gpu, The compatibility between PySCF and GPU4PySCF has been enhanced by the newly designed API in this work. Previously, all GPU4PySCF classes inherited from PySCF classes, with methods patched to use GPU functions. As more functionalities were added, the multiple inheritance design leads to many ambiguous situations when the features are defined in both PySCF and GPU4PySCF. The ambiguity increases the complexity of development dramatically. Starting from v1.0, GPU4PySCF drops the patching mechanism and explicitly redefines the same classes as in PySCF. GPU4PySCF classes and PySCF classes are constructed in parallel. A PySCF object can be converted into a GPU4PySCF object with calling to_gpu function when the functionality is implemented on GPU. The kernel execution will be accelerated with GPU. Conversely, a GPU4PySCF object can be converted back into a PySCF object using the to_cpu function, allowing PySCF-only features to be applied. Here is an example